# What Makes LLM Agent Simulations Useful for Policy? Insights From an Iterative Design Engagement in Emergency Preparedness

### v1

@@ 7.2.1. Start with Verifiable Scenarios and Build Trust Gradually
Unlike traditional tools, LLM agent simulations claim to model complex social phenomena that cannot be directly verified (Gao et al., 2023; Hou et al., 2025; Karten et al., 2025). However, our “validation filter” (Sec 6.1) finding shows that simulations only become useful when some baseline of validation against observable reality is possible. Policymakers are unlikely to accept simulation outputs they cannot check, since LLM agent simulations produce outputs that are inherently unpredictable. Instead, trust is built post-hoc, when simulated outcomes align with familiar or observable patterns.
This unpredictability is also the strength of LLM agent simulations—they allow us to explore unanticipated situations through agent interaction—but it makes the question of validation especially challenging. To address this challenge, developers and designers should first identify verifiable cases that can serve as foundations (Helbing et al., 2002) and deliberately stage this progression: begin with routine scenarios stakeholders recognize, confirm alignment with their expectations, and only then introduce speculative cases. This “trust bootstrapping” approach (Sec 6.2) allows LLM agent simulations to harness their generative capacity while extending policymakers’ confidence from validated setups to novel, unverified scenarios.
@@ 7.2.2. Use Preliminary Simulations to Elicit Tacit Knowledge
In HCI, prototyping is not just about refining and reducing errors, but about exploring ideas through progressive increases in fidelity as designers gain confidence (Zimmerman et al., 2007; Buchenau and Suri, 2000; Hartmann et al., 2006; Park et al., 2022). LLM agent simulations for policy implementation can benefit from the process in similar ways, but our “fix-it” response finding (Sec 6.3) highlights an additional benefit. With the verisimilitude of LLM-generated behaviors, even deliberately preliminary or “wrong” simulations can generate the most valuable conversations with policymakers, surfacing tacit domain knowledge that otherwise they cannot articulate upfront.
This marks a clear departure from typical AI development processes, which emphasize system accuracy and polished final outputs (Choudhuri et al., 2025; Brundage et al., 2020). All simulations are necessarily imperfect—they cannot capture an entire world. In fact, as Bonini’s paradox reminds us, models can become less useful as they attempt to capture more of a complex system (Bonini et al., 1967). Like a map that is most useful when it selectively simplifies the territory, the challenge in simulation design is not simply to increase fidelity, but to choose carefully which aspects to represent in order to make the model actionable.
Developers and designers should therefore treat interim simulations not just as flawed prototypes to be corrected, but as probes that invite critique. Framing outputs with questions such as “what’s missing here?” rather than “is this correct?” helps uncover tacit domain knowledge and institutional practices that can ultimately make the system more useful. Importantly, the missing elements are not always tied to LLM capabilities themselves. Policymakers care about the broader social context, beyond agent behavior alone. In our case, much of the feedback centered on the interaction environment—such as spatial layout, family relationships, or coordinator roles—which ultimately proved crucial for making the simulations useful for policy and its implementation (Sec 6.4). These findings underscore that the details surfaced through iterative feedback are not incidental: they are precisely what determines whether a simulation achieves the fidelity needed to inform policy implementation.
@@ 7.2.3. Set a Design Focus on Evolving Simulation Capabilities and Policy Requirements Together
User-centered design typically assumes relatively stable user needs that technology should accommodate (Abras et al., 2004; Gulliksen et al., 2003; Vredenburg et al., 2002). However, we found that simulation capabilities and policymaker requirements were inter-dependent and developed simultaneously, with neither existing in completing form at the outset (Sec 6.5). In our case, simulations began as simple demonstrations of evacuation patterns, but over time policymakers used them to experiment with coordinator layout and differentiated protocols. At the same time, several suggestions based on our simulations were not pursued—even after trust in the system was established—because of institutional constraints. One example was for local law enforcement to assume authority during bomb threats. This contrast shows that simulation and policy implementation development can only move forward together when technical possibilities and institutional boundaries are both taken into account. For complex AI systems in institutional contexts, precise design goals and useful solutions cannot be fully defined at the onset, but must be iteratively co-constructed.
Developers and designers should therefore set a design focus for evolving both simulation capabilities and policy requirements as interdependent targets, rather than treating one as subordinate to the other. This means focusing design on decision spaces where adaptation is possible—such as resource allocation, training, or communication—while building flexibility so that both the system and institutional processes can adjust as shared understanding deepens. By embedding this principle of interaction into the dual design focus from the outset, simulations can move beyond prototypes to become enduring tools for practice.
@@ 7.3. Limitations and Future Directions
Our single-domain focus limits generalizability. Emergency preparedness offers clear success metrics and observable validation opportunities that may not exist in other policy domains. The close collaborative relationships with policymakers may also have influenced outcomes in ways that differ from typical technology adoption scenarios.
@@
Ultimately, the usefulness of such tools should be judged by whether suggested changes in policy implementation actually improve safety during emergencies. Our study did not directly test implementation results. Longitudinal studies are needed to evaluate this question and to trace how simulation tools evolve over extended periods of institutional use. Future research should also test these design conditions across other policy domains and organizational contexts, and explore ways to scale iterative methods to institutions with fewer resources. Developing evaluation frameworks that center stakeholder experience, rather than traditional accuracy alone, represents another important methodological direction.
@@ 8. Conclusion
In this paper, we presented a year-long iterative design of an LLM agent simulation system with a university emergency preparedness team, showing how participatory engagement enabled simulations to move from academic demonstrations to institutionally integrated tools. Our findings demonstrate that usefulness arose not from technical fidelity alone, but from design practices that helped build trust and utility gradually. By tracing this pathway, we conclude with design implications for practitioners at the intersection of HCI, AI, and policymaking: begin with validation-feasible domains, treat imperfections as opportunities to elicit expertise, and set a design focus on simulation and policy implementation together.
As famously observed in statistics, “All models are wrong, but some are useful(Box, 1976). Our study examines how LLM agent simulations can be designed to be useful for policy, not by claiming predictive certainty, but by aligning technical possibilities with institutional needs. This work provides a stepping stone for harnessing emerging AI capabilities in service of human well-being through more effective policy implementation.
@@ A.1. Co-creating Stakeholder Maps and Process Maps
Figure A1. Developers and policymakers P1 collaboratively co-creating a stakeholder map and process map during a policy meeting. The session supported knowledge elicitation of roles, responsibilities, and workflows in emergency preparedness.:
Figure A2. Stakeholder map of emergency preparedness and response at the university, co-created with policymakers to identify actors, roles, and relationships shaping decision-making and information flows during crises.:
Figure A3. Process diagram of emergency preparedness and response at the university, documenting sequential phases from preparedness planning to active response and recovery, and highlighting interdependencies among roles.:
Figure A4. Process diagram of emergency preparedness and response tailored for severe weather scenarios, illustrating protocols for monitoring, issuing warnings, coordinating shelter, and resuming normal operations.:
Figure A5. Process diagram of emergency preparedness and response tailored for unrest or bomb threat scenarios, showing decision pathways for lockdowns, communication flows, and coordination with law enforcement.:

### v2

@@ 7.2.1. Start with Verifiable Scenarios and Build Trust Gradually
LLM agent simulations often target complex social phenomena that cannot be directly verified (Gao et al., 2023; Hou et al., 2025; Karten et al., 2025). Our “validation filter” finding (Sec 6.1) shows that, in practice simulations only become policymakers could first validate outputs against observable or familiar scenarios. Because LLM-generated behaviors are inherently stochastic, trust was not established through formal evaluation alone, but through post-hoc alignment with known patterns of behavior. This unpredictability is also the strength of LLM agent simulations—they allow us to explore unanticipated situations through agent interaction—but it makes the question of validation especially challenging.
Developers and designers should therefore treat validation not as a final evaluation step, but as an early design constraint that shapes which scenarios are explored first. Starting from routine, verifiable cases allows simulations to establish baseline credibility, which can then be extended to more speculative or unobservable situations. This staged progression — what we describe as “trust bootstrapping” (Sec 6.2) — enables LLM agent simulations to leverage their generative capacity while remaining usable for policy practice.
@@ 7.2.2. Use Preliminary Simulations to Elicit Tacit Knowledge
In HCI, prototyping is not just about refinement, but about exploration through progressively increasing fidelity (Zimmerman et al., 2007; Buchenau and Suri, 2000; Hartmann et al., 2006; Park et al., 2022). LLM agent simulations can serve as a similar role in policy contexts; however, our “fix-it” response finding (Sec 6.3) highlights an additional benefit. Because LLM-generated behaviors appear socially plausible, even deliberately preliminary or “wrong” simulations can provoke rich critique, surfacing tacit domain knowledge that policymakers often cannot articulate in advance.
This contrasts with typical AI development practices that emphasize accuracy and polished outputs (Choudhuri et al., 2025; Brundage et al., 2020). All simulations are necessarily imperfect—they cannot capture an entire world. As Bonini’s paradox reminds us, increasing model complexity does not necessarily increase usefulness (Bonini and others, 1967). Like a map that is most useful when it selectively simplifies the territory, the challenge in simulation design is not simply to increase fidelity, but to choose carefully which aspects to represent in order to make the model actionable.
Developers should therefore treat interim simulations not simply as flawed prototypes to be corrected, but as probes that invite critique. Framing outputs with questions such as “what’s missing here?” rather than “is this correct?” helps uncover tacit domain knowledge and institutional practices that can ultimately make the system more useful. Importantly, the missing elements are not always tied to LLM capabilities themselves, but often to how the broader interaction context is represented. Policymakers care about the broader social context, beyond agent behavior alone. In our case, much of this feedback concerned the interaction environment—such as spatial layout, family relationships, or coordinator roles—which ultimately proved central to making the simulations useful for policy implementation (Sec 6.4). These findings underscore that the details surfaced through iterative feedback are not incidental: they are precisely what determines whether a simulation achieves the fidelity needed to inform policy implementation. At the same time, the “fix-it” dynamic can become counterproductive when early simulations generate an unbounded list of realism demands. In our study, iteration remained productive when critiques could be translated into concrete representational changes and linked to decisions that policymakers could plausibly act on.
@@ 7.2.3. Set a Design Focus on Evolving Simulation Capabilities and Policy Requirements Together
User-centered design often assumes relatively stable user needs that technology should accommodate (Abras et al., 2004; Gulliksen et al., 2003; Vredenburg et al., 2002). In contrast, we found that simulation capabilities and policymaker requirements co-evolved, with neither fully specified at the outset (Sec 6.5). In our case, simulations began as simple demonstrations of evacuation patterns, but over time policymakers used them to explore coordinator placement and differentiated protocols. At the same time, some simulation-informed ideas were not pursued due to institutional constraints (e.g., law enforcement authority during bomb threats), highlighting that technical possibilities and organizational boundaries must be considered together. More broadly, simulation insights were easiest to adopt when the same team that interpreted the results could also change the relevant procedures. Adoption became more difficult when changes required coordination across organizational boundaries or shifts in incident command.
For complex AI systems in institutional contexts, useful design goals cannot be fixed in advance but must be iteratively co-constructed. Developers should therefore focus on decision spaces where adaptation is possible—such as training, resource allocation, or communication—while maintaining flexibility for both simulation capabilities and institutional processes to evolve. This co-evolution enables simulations to move beyond isolated demonstrations and become sustained tools for policy implementation.
@@ 7.3. Limitations and Future Directions
Our single-domain focus limits generalizability. Emergency preparedness offers clear success metrics and observable validation opportunities that may not exist in other policy domains. In addition, our strongest validation opportunities involved routine, non-emergency conditions. We did not have the chance to observe real-world behavior under active threat, where fear, uncertainty, and authority dynamics may alter coordination in more fundamental ways. The close collaborative relationships with policymakers may also have influenced outcomes in ways that differ from typical technology adoption scenarios.
@@
Ultimately, the usefulness of such tools should be judged by whether suggested changes in policy implementation actually improve safety during emergencies. Our study did not directly test implementation results. Longitudinal studies are needed to evaluate this question and to trace how simulation tools evolve over extended periods of institutional use. A complementary near-term direction is to examine intermediate outcomes that precede rare emergencies, such as whether simulation-based training measurably changes coordinator behavior during drills or improves routine crowd management. Future research should also test these design conditions across other policy domains and organizational contexts, and explore ways to scale iterative methods to institutions with fewer resources. Developing evaluation frameworks that center stakeholder experience, rather than traditional accuracy alone, represents another important methodological direction.
Another important direction for future work is comparative evaluation with traditional agent-based models. LLM agent simulations are a form of agent-based modeling, but differ from traditional approach in how agent behavior is specified: rather than relying on hand-crafted rules or physics-driven dynamics alone, LLM agent simulations use language models to generate semantically rich, context-sensitive behavior. These differences affect how simulations are designed, validated, and integrated into policy practice. Accordingly, meaningful evaluation in policy contexts would examine how LLM-based approaches complement — and trade off with — the established strengths of traditional agent-based models across dimensions such as interpretability, trust calibration, stakeholder engagement, and alignment with institutional constraints (Lempert, 2002).
@@ 8. Conclusion
In this paper, we presented a year-long iterative design engagement with a university emergency preparedness team to examine how LLM agent simulations can become useful for policy practice. Rather than evaluating simulations as predictive or optimization tools, we traced how participatory design enabled them to evolve from academic demonstrations into institutionally integrated resources for planning, training, and reflection. Our findings show that usefulness did not arise from technical fidelity alone, but from design practices that gradually built trust, relevance, and shared understanding between developers and policymakers.
By tracing this pathway, we articulated three design process implications for practitioners working at the intersection of HCI, AI, and policymaking: begin with scenarios that admit some form of validation, treat imperfections as opportunities to surface tacit expertise, and co-evolve simulation capabilities with policy implementation needs. Together, these implications shift attention from asking whether LLM agent simulations are “accurate enough” toward understanding how they can be responsibly embedded in real organizational workflows.
As famously observed in statistics, “All models are wrong, but some are useful(Box, 1976). Our contribution is to show how LLM agent simulations can be designed to be useful for policy practice, not by claiming predictive certainty, but by aligning technical possibilities with institutional logics, constraints, and decision-making processes. More broadly, this work suggests that the value of emerging AI systems in high-stakes domains may hinge less on advances in model capability than on the design processes that connect those capabilities to human judgment, institutional practice, and organizational sensemaking. We hope this study provides a foundation for future HCI research on designing AI systems that support—not replace—policy practice. We also hope it informs more careful, grounded integration of LLM-based simulations into real-world institutions.
@@ A.1. Co-creating Stakeholder Maps and Process Maps
Figure A1. Developers and policymakers P1 collaboratively co-creating a stakeholder map and process map during a policy meeting. The session supported knowledge elicitation of roles, responsibilities, and workflows in emergency preparedness.
A photo of two people seated at a table in a meeting room, working together on a white board. They are sketching maps with markers to document stakeholders and processes for emergency preparedness.
Figure A2. Stakeholder map of emergency preparedness and response at the university, co-created with policymakers to identify actors, roles, and relationships shaping decision-making and information flows during crises.
A network diagram showing university units, local government offices, emergency responders, and student groups, with labeled connections illustrating communication and authority relationships in emergency preparedness.
Figure A3. Process diagram of emergency preparedness and response at the university, documenting sequential phases from preparedness planning to active response and recovery, and highlighting interdependencies among roles.
A flowchart with boxes and arrows showing stages of preparedness (planning, drills), active response (alerts, evacuation), and recovery (debrief, after-action reporting), with actors assigned at each stage.
Figure A4. Process diagram of emergency preparedness and response tailored for severe weather scenarios, illustrating protocols for monitoring, issuing warnings, coordinating shelter, and resuming normal operations.
A process map with steps beginning at weather monitoring, progressing through warning issuance, shelter coordination, and eventual campus reopening, with designated responsibilities for staff and emergency services.
Figure A5. Process diagram of emergency preparedness and response tailored for unrest or bomb threat scenarios, showing decision pathways for lockdowns, communication flows, and coordination with law enforcement.
A flowchart with branching paths showing actions under unrest or bomb threat: initiating lockdowns, coordinating with police, informing students and staff, and transitioning from response to resolution.